# Deep Learning Atmospheric Models Reliably Simulate Out-of-Sample Land Heat and Cold Wave Frequencies

### v1

@@ 4 Conclusion and Discussion
We have evaluated the ability of deep-learning (DL) general circulation models (NGCM and DLESyM) to simulate land heatwaves and coldwaves during 1900–2010, comparing their performance with a physical model (HiRAM) using simulations following an AMIP protocol. Despite being trained on ERA5 data from 1980–2020, both DL models successfully reproduce the frequency and spatial patterns of heatwave and coldwave events during the out-of-sample period (1900–1960), with skill comparable to HiRAM, and mostly high correlations with reanalysis verification (20CRv3). Exceptions include the tropics, where we find that the models have higher autocorrelation that 20CRv3, and North Asia and central North America, which we speculate may be due to changes in land-surface conditions not present in any of the models.
Model architecture appears to affect temperature extremes frequency through the autocorrelation of the surface temperature field. The fully data-driven model DLESyM tends to overestimate extreme event frequency and has the highest temporal autocorrelation, while the physics–DL hybrid NGCM produces event frequencies more aligned with HiRAM, and correspondingly similar autocorrelation. The higher temperature autocorrelation in fully DL-based models may result from the absence of explicit physical constraints and a tendency to produce smoother, more persistent temperature anomalies due to learning objectives—such as loss functions (e.g., mean squared error)—that favor minimizing short-term variability. Additionally, the large ensemble sizes enabled by the efficiency of DL models improve the stability and robustness of their climatological estimates, particularly compared to HiRAM’s limited 5-member ensemble.

### v2

@@ 4 Conclusion and Discussion
We have evaluated the ability of deep-learning (DL) general circulation models (NGCM and DLESyM) to simulate land heatwaves and coldwaves during 1900–2010, comparing their performance with a physical model (HiRAM) using simulations following an AMIP protocol. Despite being trained on ERA5 data from 1980–2020, both DL models successfully reproduce the frequency and spatial patterns of heatwave and coldwave events during the out-of-sample period (1900–1960), with skill comparable to HiRAM, and mostly high correlations with reanalysis verification (20CRv3). Exceptions include the tropics, where we find that the models have higher autocorrelation that 20CRv3, and North Asia and central North America, which we speculate may be due to changes in land-surface conditions not present in any of the models. Furthermore, by directly comparing temperature distributions across six cities, we show that the DL models are able to capture the tails of the temperature distributions (temperature extremes).
The degree of physical components appears to affect temperature extremes frequency through the autocorrelation of the surface temperature field. The fully data-driven model DLESyM tends to overestimate extreme event frequency and has the highest temporal autocorrelation, while the physics–DL hybrid NGCM produces event frequencies more aligned with HiRAM, and correspondingly similar autocorrelation. The higher temperature autocorrelation in fully DL-based models (DLESyM and ACE2-ERA5) may result from the absence of explicit physical constraints and a tendency to produce smoother, more persistent temperature anomalies due to learning objectives—such as loss functions (e.g., mean squared error)—that favor minimizing short-term variability. Additionally, the large ensemble sizes enabled by the efficiency of DL models improve the stability and robustness of their climatological estimates, particularly compared to HiRAM’s limited 5-member ensemble.
Figure 5: Distribution of temperature anomalies from 1900–1960. Log-scaled temperature probability density distributions for DJF (left) and JJA (right) at Helsinki, Paris, Washington DC, Cairo, Mexico City, and Sydney. Anomalies are computed relative to the 1980–2010 reference period. Black denotes Berkeley Earth (BE), red denotes NGCM, orange denotes DLESyM, purple denotes 20CR, and green denotes HiRAM; lines are distributed above and below the abscissa for clarity. Shading represents the distribution, with the thin vertical solid line indicating the median (50th percentile) and dashed lines indicating the 1st and 99th percentiles.
@@ Acknowledgments
Supporting Information for “Evaluation of Out-of-Sample Land Heat and Cold Waves in Deep Learning-Based Atmospheric Models"
Contents of this file
  1. 1.
    Figures S1 to S12
  2. 2.
    Table S1
Table 1: Summary of Models and Experimental Setup.
Attribute NGCM DLESyM HiRAM
Model Type Physics–DL Hybrid Purely Data-Driven Physical GCM
Horizontal Resolution $\sim$ 2.8° $\sim$ $\sim$50 km (C180)
Boundary Forcing SST & SIC SST SST & SIC
Training Period 1979–2017 (ERA5) 1983–2016 (ERA5/ISCCP) N/A (Physics-based)
Ensemble Size 100 100 5
@@
Figure S12: Link between SST analogue similarity and extremes frequency similarity. For each out-of-sample year (1900–1960), the in-sample year (1979–2017) with the highest SST anomaly pattern correlation is identified as its analogue year. Scatter plots show the relationship between SST anomaly similarity (correlation, x-axis) and the model-simulated frequency pattern similarity of heatwaves (left panels) and coldwaves (right panels) for DLESYM, HiRAM, and NGCM. Red lines indicate linear regression fits, with panel titles reporting correlation coefficients and $p$-values. Panel labels (a–f) correspond to each model–extreme combination.